OpenStudy (anonymous):

solution? Square root(2x+13)-5=x

OpenStudy (anonymous):

remove the square root to make things easier

OpenStudy (mathmale):

We have to get rid of the square root operator. Squaring a square root will do that. However, squaring the left side of your equation will make more work, not less work. So, think about what you might to to that equation BEFORE you square both sides.

OpenStudy (mathmale):

sorry ... "think about what you might DO to that equation" is what I meant.

OpenStudy (anonymous):

the only thing i can think to do first would be to move the -5 over to the other side of the equation.

OpenStudy (anonymous):

@mathmale is that what you mean?

OpenStudy (mathmale):

Yes, exactly. Just add 5 to both sides of the equation to accomplish that.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then what

OpenStudy (mathmale):

Now your left side consists solely of the square root of some quantity, and the right side of a simple binomial. You can now go ahead and square both sides. This will remove the Sqrt from the left side.

OpenStudy (anonymous):

ok so it would look like 2x+13=(x+5)^2

OpenStudy (anonymous):

@mathmale ok well if that is correct then can you help me with my next problem? which is ( square root (250x^16) ) / (square root (2x) )

OpenStudy (mathmale):

Morgan-Alana: What you have given me in your second-from-last response is correct, but I'm afraid you're not off the hook yet! ;) You still have to expand (x+5)^2 and solve the resulting equation for x. Note: You'll get more than one possible solution; you'll need to check both in the original equation and determine which, if any, is actually a solution .

OpenStudy (anonymous):

@mathmale ok then from there i got 2x+13=x^2+25 then i moved the x^2 over to the left side and the +13 over to the right side and then that gave me 2x-x^2=12 from there i am lost

OpenStudy (amoodarya):

\[\sqrt{2x+13}-5 =x \] \[\sqrt{2x+13} =x +5\] \[(\sqrt{2x+13} =x +5)^2\]

OpenStudy (amoodarya):